OpenStudy (anonymous):

Solve for x. - 2x + 5y = - 8 6x - 15y = 24 A. x is undefined. B. x is any real number. C. x = 0 D. x = - 3

OpenStudy (arabpride):

Try to set it up in 'y=mx+b' form

OpenStudy (arabpride):

- 2x + 5y = - 8 +2x +2x 5y = 2x - 8 /5 /5 y = 2/5x -8/5 ~~~~~~~~~~~~~~~~~~~ 6x - 15y = 24 -6x -6x -15y = -6x + 24 /-15 /-15 y = 6/15x -24/15 ~~~~~~~~~~~~~~~~~~~~ >>>>y = 2/5x -8/5<<<< and >>>>y = 6/15x -24/15<<<<